# Data-Driven Prescriptive Analytics Applications: A Comprehensive Survey

### v2

@@ Abstract
Prescriptive Analytics (PSA), an emerging business analytics field suggesting concrete options for solving business problems, has seen an increasing amount of interest after more than a decade of multidisciplinary research. This paper is a comprehensive survey of existing applications within PSA in terms of their use cases, methodologies, and possible future research directions. To ensure a manageable scope, we focus on PSA applications that develop data-driven, automatic workflows, i.e., Data-Driven PSA (DPSA). Following a systematic methodology, we identify and include 104 papers in our survey. As our key contributions, we derive a number of novel taxonomies of the field and use them to analyse the field’s temporal development. In terms of use cases, we derive 10 application domains for DPSA, from Healthcare to Manufacturing, and subsumed problem types within each. In terms of individual method usage, we derive 5 method types and map them to a comprehensive taxonomy of method usage within DPSA applications, covering mathematical optimization, data mining and machine learning, probabilistic modelling, domain expertise, as well as simulations. As for combined method usage, we provide a statistical overview of how different method usage combinations are distributed and derive 2 generic workflow patterns along with subsumed workflow patterns, combining methods by either sequential or simultaneous relationships. Finally, we derive 4 possible research directions based on frequently recurring issues among surveyed papers, suggesting new frontiers in terms of methods, tools, and use cases.
@@ 1 Introduction
    Finally, we derive 4 possible research directions reflecting the current state of the art, based on an analysis of frequently recurring issues and temporal trends within DPSA applications.
The remainder of this paper is structured as follows: First, we provide a systematic analysis of how the present survey marks a relevant scientific contribution compared to existing PSA surveys in Section 2. Section 3 covers relevant background for readers unfamiliar with PSA. Then Section 4 provides a definition of DPSA, to delimit the scope of this survey. Sections 5-7 move on to present the chosen survey methodology, choices made as part of the survey process, along with a statistical overview of included papers, respectively. Section 8, Sections 9-11, and Section 12 address each research question in turn, providing a comprehensive framework of problem and solution types found within existing application research, along with a number of possible future research directions. We provide concluding remarks in Section 13.
@@ 2 Comparison to existing PSA surveys
First, we note that a number of existing surveys don’t cover any papers from this decade [13, 10, 9, 14]. As our analysis in Section 7 shows, 76.9% of our surveyed papers were published within the 2020-2024 year range, with the number of papers having accelerated significantly over the past three years. Looking solely at the recency of content, this survey therefore provides substantially updated knowledge.
@@ 3 Background on PSA
PSA arose as a term within the field of business analytics in the previous decade [5]. A broad, yet comprehensive definition of business analytics, derived from a survey of existing definitions within the field [3,  p. 138], goes as follows:
@@ 4 PSA definitions and survey scope
To ensure a well-defined scope, it was ultimately deemed necessary to use a more specific version of Definition 2 in this survey. Throughout history, many prescriptive approaches have been used within human organizational endeavours: from ancient astronomy using the constellations to time when to harvest the crops [34,  p. 4], to modern operations research arriving to support military efforts during the Second World War [33]. Even disregarding definitional issues, a single survey paper cannot realistically aspire to cover everything written within the scope of Definition 2 since the dawn of history, let alone entire fields like operations research, which came to be more than 80 years ago [35], with a corresponding number of papers to boot.
@@ 5 Survey methodology
This section outlines the utilized survey methodology, taking its onset in the recommendations followed by previous surveys within PSA [23, 9]. Sections 6 and 7 provides a statistical summary of the search process and identified papers, respectively. Throughout, we follow the PRISMA 2020 guidelines for systematic literature review reporting [41].
@@ 8 Problem types
Developing the aforementioned categories was done iteratively during the full text screening phase (cf. Section 6), using previous efforts as the initial expectation (e.g., [14, 9]), yet ultimately ending up with an original taxonomy that would fit the set of surveyed papers - merging, splitting, deleting and inventing categories as necessary. Specifically, the heuristic followed was that a category should contain at least three papers, in order to avoid an overly baroque set of categories, yet at most 20%, or 20 papers, in order for the categories to work as a more useful index within the context of this survey.
@@
The sheer variety of problem types covered by Table 2 suggests that a broad notion of ”business problem solving” (cf. Definition 1) is reflected by the literature as a whole, concerning organizational value creation [154] for many different kinds of stakeholders, be it in hospitals, factories, universities, social services, grocery stores, or elsewhere. Still, due to their common core methodology (cf. Definition 3), patterns in terms of choice of methods and workflows can be identified across use cases and application domains, as Sections 9-11 explore further.
@@ 9 Method types
This section addresses part of RQ2 by summarizing the array of individual methods used for DPSA in the set of surveyed papers. Later on, Section 10 addresses RQ2 in terms of how these methods are combined in different workflow patterns. In both Sections 9 and 10, there is a focus on offering new taxonomies and conceptual models to form a structured understanding of existing DPSA methodologies.
@@ 9.1 Data Mining and Machine Learning (DM/ML)
This method type covers statistical learning and knowledge discovery from historical datasets [155,  p. 5]. As shown in Figure 7, predictive DM/ML encompasses a large array of methods. Indeed, one might find that this subtree reads like a data mining or ML textbook, with comprehensive coverage of most popular algorithms and approaches.
@@ 9.3 Probabilistic Modelling (PROB)
This method type covers techniques modelling uncertain causal relationships. While some prescriptive methods already discussed, such as Stochastic Programming and Reinforcement Learning (cf. Sections 9.2 and 9.1) overlap with PROB conceptually, this category is ultimately more closely associated with a set of predictive methods (cf. Figure 7).
@@ 10.4 Predicting-While-Prescribing (PWP) workflows
In a sense, Reinforcement Learning embodies the PWP pattern. Some reinforcement learning algorithms solve predictive and prescriptive problems concurrently - the value and policy function, respectively estimating the prospect of being in a state and the optimal way to navigate from it, being ”moving targets” for each other until convergence [168, p. 97]. Whether they use data-driven reinforcement learning as a stand-alone approach [58, 68, 67, 116, 50] or supplement it with additional predictive methods used in an online fashion [125, 123, 71], papers using reinforcement learning all at least embody the PWP pattern in such a sense.
@@ 11 Synthesis on methods
As Sections 9 and 10 have covered while addressing RQ2, several definite methodological trends characterize the state of current DPSA application research, both in terms of individual methods and their combinations.
DPSA applications can be said to pivot around five method types, or ”clusters” if such imagery is preferred, among which DM/ML and OPT dominate (cf. Figure 5), and has become increasingly dominant in recent years (cf. Figure 6), yet with several viable alternatives in the shape of PROB, EXP, and SIM. While a pronounced emphasis on tree-based ensembles for predictions and Linear Programming methods for prescriptions was found (cf. Sections 9.1 and 9.2), individual method choices varied widely between applications, exhibiting a textbook level of coverage for the most popular method types. To our knowledge, the various prescriptive EXP methods described in Section 9.4 have not been covered explicitly elsewhere, despite the popularity of prescriptive EXP (cf. Figure 9).
@@ 12 Promising research directions
Based on pervasive patterns found in the set of surveyed papers, we derive four promising research directions (RD1-4) to address RQ3, presented in the following subsections.
12.1 The promise of Big DPSA
A few papers describe their own work as involving Big Data or evidently use distributed processing to handle large datasets [92, 124, 144, 80]. Yet these papers tend to omit implementation details, in some cases possibly for confidential reasons [144], while others only analyse single-node-sized datasets [80].
Echoing previous surveys on PSA [9, 10], we thus still find that the promise of Big Data has not come to fruition within DPSA application research, and propose as a possible research direction to:
RD1: Develop novel methods and tools that support DPSA applications on Big Data.
@@ 12.2 Under-explored application domains
As found in connection to RQ1 (cf. Section 8), DPSA has already found use in a variety of identified application domains, and our temporal analysis demonstrates an ebb and flow in terms of relative prominence among these domains over time (cf. Figure 4). Still, several prominent real-world business domains have been left (almost) unexplored by the set of surveyed papers as a whole. Using UN’s International Standard Industrial Classification (ISIC), revision 4, [172], we have identified eleven business domains in which at most three DPSA applications were found in this survey. Leaving out the miscellaneous sections S, T, and U [172,  p. 43], the eight remaining business domains and section letters, along with associated surveyed papers, if any, are the following:
@@
To further the proliferation of DPSA applications towards new frontiers, cross-fertilizing an already interdisciplinary field with new problems and solutions, we therefore propose as a possible research direction to:
RD2: Develop impactful DPSA applications within under-explored business domains.
@@ 12.3 Methodological alternatives to (mixed-)integer linear programming
RD3: Develop scalable, expressive, usable, and reliable prescriptive DM/ML methods for DPSA applications.
12.4 Flexible tools for DPSA
@@
From our analyses concerning RQ2 (cf. Sections 9 and 10), we note that DPSA applications display a high degree of methodological diversity, both in terms of individual methods and how they are combined, having diversified further over time (cf. Figure 10). Accommodating such a wide variety of workflows poses a challenge to developers of dedicated DPSA tools: As is evident from the aforementioned sections, a PTP pattern with ML for predictions and Linear Programming for prescriptions is but one way to structure a DPSA workflow among many. As for (M)ILP specifically, we furthermore saw in Section 12.3 that significant workflow variations even exist within this one category.
@@
Dedicated DPSA tools, integrating predictive and prescriptive technologies with a data management layer, might nonetheless help their users avoid improvised ”duct-tape-style” workflows that are both error-prone to develop and exhibit poor performance [10]. Solutions like SolveDB+, a relational database management system for PSA, have already been developed in this direction [8, 49], yet they only support a particular subset of workflows discussed in this survey.
@@
How to best offer tool support for DPSA given its methodological diversity, be it in one unified tool, in several tools with good mutual integration, or otherwise, we leave as a possible research direction, more formally to:
@@
RD4: Develop flexible and efficient tools for common DPSA workflow patterns.
@@ 13 Conclusion
In answering our three research questions in Sections 8-12, we have provided new conceptual models and temporal analyses of the field as a whole, both in terms of problem types (cf. Section 8), method types (cf. Sections 9), and workflow patterns (cf. Section 10). Based on temporal trends and pervasive Gordian Knots found within the set of surveyed papers, we have furthermore provided a foundation for future work by proposing a number of possible research directions (cf. RD1-4 from Section 12), both in terms of new methods, tools, and use cases. We can mnemonically summarize the key contributions in our conceptual framework as 10 applications domains, 5 method types, 2 generic workflow patterns, and 4 research directions. These contributions form a comprehensive model of the current state of the DPSA application research field and where it might be going in the future.

### v3

@@ Abstract
Prescriptive Analytics (PSA), an emerging business analytics field suggesting concrete options for solving business problems, has seen an increasing amount of interest after more than a decade of multidisciplinary research. This paper is a comprehensive survey of existing applications within PSA in terms of their use cases, methodologies, and possible future research directions. To ensure a manageable scope, we focus on PSA applications that develop data-driven, automatic workflows, i.e., Data-Driven PSA (DPSA). Following a systematic methodology, we identify and include 104 papers in our survey. As our key contributions, we derive a number of novel taxonomies of the field and use them to analyse the field’s temporal development. In terms of use cases, we derive 10 application domains for DPSA, from Healthcare to Manufacturing, and subsumed problem types within each. In terms of individual method usage, we derive 5 method types and map them to a comprehensive taxonomy of method usage within DPSA applications, covering mathematical optimization, data mining and machine learning, probabilistic modelling, domain expertise, as well as simulations. As for combined method usage, we provide a statistical overview of how different method usage combinations are distributed and derive 2 generic workflow patterns along with subsumed workflow patterns, combining methods by either sequential or simultaneous relationships. Finally, we derive 5 possible research directions based on frequently recurring issues among surveyed papers, suggesting new frontiers in terms of methods, tools, and use cases.
@@ 1 Introduction
    Finally, we derive 5 possible research directions reflecting the current state of the art, based on an analysis of frequently recurring issues and temporal trends within DPSA applications.
The remainder of this paper is structured as follows: First, we provide a systematic analysis of how the present survey marks a relevant scientific contribution compared to existing PSA surveys in Section 2. Section 3 covers relevant background for readers unfamiliar with PSA. Then Section 4 provides a definition of DPSA, to delimit the scope of this survey. Section 5 moves on to present the chosen survey methodology, choices made as part of the survey process, along with a statistical overview of included papers. Section 6, Sections 7-9, and Section 10 address each research question in turn, providing a comprehensive framework of problem and solution types found within existing application research, along with a number of possible future research directions. We provide concluding remarks in Section 11.
@@ 2 Comparison to existing PSA surveys
First, we note that a number of existing surveys don’t cover any papers from this decade [13, 10, 9, 14]. As our analysis in Section 5.5 shows, 76.9% of our surveyed papers were published within the 2020-2024 year range, with the number of papers having accelerated significantly over the past three years. Looking solely at the recency of content, this survey therefore provides substantially updated knowledge.
@@ 3 Background on PSA
PSA arose as a term within the field of business analytics in the previous decade [5]. A broad, yet comprehensive definition of business analytics, derived from a survey of existing definitions within the field [3], goes as follows:
@@ 4 PSA definitions and survey scope
To ensure a well-defined scope, it was ultimately deemed necessary to use a more specific version of Definition 2 in this survey. Throughout history, many prescriptive approaches have been used within human organizational endeavours: from ancient astronomy using the constellations to time when to harvest the crops [34], to modern operations research arriving to support military efforts during the Second World War [33]. Even disregarding definitional issues, a single survey paper cannot realistically aspire to cover everything written within the scope of Definition 2 since the dawn of history, let alone entire fields like operations research, which came to be more than 80 years ago [35], with a corresponding number of papers to boot.
@@ 5 Survey methodology
Sections 5.1-5.3 outline the utilized survey methodology, taking their onset in recommendations followed by previous surveys within PSA [23, 9]. Sections 5.4 and 5.5 provide a statistical summary of the search process and identified papers, respectively. Throughout, we follow the PRISMA 2020 guidelines for systematic literature review reporting [41].
@@ 6 Problem types
Developing the aforementioned categories was done iteratively during the full text screening phase (cf. Section 5.4), using previous efforts as the initial expectation (e.g., [14, 9]), yet ultimately ending up with an original taxonomy that would fit the set of surveyed papers - merging, splitting, deleting and inventing categories as necessary. Specifically, the heuristic followed was that a category should contain at least three papers, in order to avoid an overly baroque set of categories, yet at most 20%, or 20 papers, in order for the categories to work as a more useful index within the context of this survey.
@@
The sheer variety of problem types covered by Table 2 suggests that a broad notion of ”business problem solving” (cf. Definition 1) is reflected by the literature as a whole, concerning organizational value creation [154] for many different kinds of stakeholders, be it in hospitals, factories, universities, social services, grocery stores, or elsewhere. Still, due to their common core methodology (cf. Definition 3), patterns in terms of choice of methods and workflows can be identified across use cases and application domains, as Sections 7-9 explore further.
@@ 7 Method types
This section addresses part of RQ2 by summarizing the array of individual methods used for DPSA in the set of surveyed papers. Later on, Section 8 addresses RQ2 in terms of how these methods are combined in different workflow patterns. In both Sections 7 and 8, there is a focus on offering new taxonomies and conceptual models to form a structured understanding of existing DPSA methodologies.
@@ 7.1 Data Mining and Machine Learning (DM/ML)
This method type covers statistical learning and knowledge discovery from historical datasets [155]. As shown in Figure 7, predictive DM/ML encompasses a large array of methods. Indeed, one might find that this subtree reads like a data mining or ML textbook, with comprehensive coverage of most popular algorithms and approaches.
@@ 7.3 Probabilistic Modelling (PROB)
This method type covers techniques modelling uncertain causal relationships. While some prescriptive methods already discussed, such as Stochastic Programming and Reinforcement Learning (cf. Sections 7.2 and 7.1) overlap with PROB conceptually, this category is ultimately more closely associated with a set of predictive methods (cf. Figure 7).
@@ 8.4 Predicting-While-Prescribing (PWP) workflows
In a sense, Reinforcement Learning embodies the PWP pattern. Some reinforcement learning algorithms solve predictive and prescriptive problems concurrently - the value and policy function, respectively estimating the prospect of being in a state and the optimal way to navigate from it, being ”moving targets” for each other until convergence [168]. Whether they use data-driven reinforcement learning as a stand-alone approach [58, 68, 67, 116, 50] or supplement it with additional predictive methods used in an online fashion [125, 123, 71], papers using reinforcement learning all at least embody the PWP pattern in such a sense.
@@ 9 Synthesis on methods
As Sections 7 and 8 have covered while addressing RQ2, several definite methodological trends characterize the state of current DPSA application research, both in terms of individual methods and their combinations.
DPSA applications can be said to pivot around five method types, or ”clusters” if such imagery is preferred, among which DM/ML and OPT dominate (cf. Figure 5), and has become increasingly dominant in recent years (cf. Figure 6), yet with several viable alternatives in the shape of PROB, EXP, and SIM. While a pronounced emphasis on tree-based ensembles for predictions and Linear Programming methods for prescriptions was found (cf. Sections 7.1 and 7.2), individual method choices varied widely between applications, exhibiting a textbook level of coverage for the most popular method types. To our knowledge, the various prescriptive EXP methods described in Section 7.4 have not been covered explicitly elsewhere, despite the popularity of prescriptive EXP (cf. Figure 9).
@@ 10 Promising research directions
Based on pervasive patterns found in the set of surveyed papers, we derive five promising research directions (RD1-5) to address RQ3, presented in the following subsections.
@@ 10.1 Under-explored application domains
As found in connection to RQ1 (cf. Section 6), DPSA has already found use in a variety of identified application domains, and our temporal analysis demonstrates an ebb and flow in terms of relative prominence among these domains over time (cf. Figure 4). Still, several prominent real-world business domains have been left (almost) unexplored by the set of surveyed papers as a whole. Using UN’s International Standard Industrial Classification (ISIC), revision 4, [169], we have identified eleven business domains in which at most three DPSA applications were found in this survey. Leaving out the miscellaneous sections S, T, and U [169], the eight remaining business domains and section letters, along with associated surveyed papers, if any, are the following:
@@
Several reasons could underlie the sparsity of DPSA applications within these specific areas. For traditionally asset-heavy, field-based industries like agriculture (A), mining (B), water/waste (E), and construction (F), data collection, e.g., through sensors, might still generally be too fragmented, manual, or sparse to properly support DPSA workflows. Within construction (F), arts and entertainment (R), along with real estate (L), decision processes traditionally being highly non-routine, subjective, or context-dependent, and thus harder to generalize, might have hindered adoption. Within real estate (L) along with finance and insurance (K), adoption of DPSA might be hindered by highly regulated, high-stakes decision processes with a need for higher auditability and explainability than current DPSA solutions support. For accommodation and food services (I), arts and entertainment (R), along with agriculture (A), the lack of adoption might be a matter of resource-efficiency: These industries might contain many small and medium-sized enterprises who lack the necessary data basis, IT infrastructure, employee skills, etc. to adopt current DPSA solutions. Note that all of the above highlighted issues are merely hypothetical possibilities. The real-world challenges of implementing DPSA within these domains still remain to be uncovered and solved. This is an interesting direction for future research in itself.
To further the proliferation of DPSA applications towards new frontiers, cross-fertilizing an already interdisciplinary field with new problems and solutions, we thus propose as a possible research direction to:
RD1: Develop impactful DPSA applications within under-explored business domains.
@@ 10.2 Methodological alternatives to (mixed-)integer linear programming
RD2: Develop scalable, expressive, usable, and reliable prescriptive DM/ML methods for DPSA applications.
10.3 Hybrid processing solutions for Big DPSA
A few surveyed papers describe their own work as involving Big Data or evidently use distributed processing to handle large datasets [92, 124, 144, 80]. Yet these papers tend to omit implementation details, in some cases possibly for confidential reasons [144], while others only analyse single-node-sized datasets [80].
The promise of Big Data being a loud refrain inside PSA narratives while only being a faint whimper inside actual DPSA applications is surprising on the surface. We shall not speculate on human motives and ascribe this finding to, e.g., hypocritical marketing or a lack of industrial buy-in. We rather hypothesize that there might be a simple, technical reason behind this situation: With current methods, DPSA is actually predominantly compute-intensive. As Section 10.2 discusses, applications currently struggle with the scalability of popular prescriptive methods like (M)ILP with respect to the number of decision variables and constraints - becoming a computational bottleneck way before any Big Data issues attain relevance.
Like previous surveys on PSA [9, 10], we still find that the promise of Big Data has not come to fruition within DPSA application research. Unlike previous works, we however propose that unlocking the Value of Big DPSA requires hybrid processing, handling both data- and compute-intensiveness. Creating such solutions introduces compounded scalability challenges. Namely, it requires integrating high-throughput data pipelines with complex optimization and search tasks that are notoriously difficult to parallelize efficiently - each with their own compute profiles and fault tolerance mechanisms [174, 37]. In sum, we propose as a research direction to:
RD3: Develop novel architectures, methods, and tools that support hybrid processing of Big DPSA workflows.
10.4 Dedicated tools for emerging DPSA workflow patterns
@@
From our analyses concerning RQ2 (cf. Sections 7 and 8), we note that DPSA applications display a high degree of methodological diversity, both in terms of individual methods and how they are combined, having diversified further over time (cf. Figure 10). Accommodating such a wide variety of workflows poses a challenge to developers of dedicated DPSA tools: As is evident from the aforementioned sections, a PTP pattern with ML for predictions and Linear Programming for prescriptions is but one way to structure a DPSA workflow among many. As for (M)ILP specifically, we furthermore saw in Section 10.2 that significant workflow variations even exist within this one category.
@@
Dedicated DPSA tools, integrating predictive and prescriptive technologies with a data management layer, might nonetheless help their users avoid improvised ”duct-tape-style” workflows that are both error-prone to develop and exhibit poor performance [10]. Solutions like SolveDB+, a relational database management system for PSA, have already been developed in this direction [8, 49], yet they only target a particular subset of workflows discussed in this survey, namely the ones combining predictive DM/ML with prescriptive OPT.
As seen in Figure 10, applications relying on prescriptive DM/ML or EXP in their workflows have been on the rise in recent years. We therefore expect a correspondingly growing tool gap within this niche. As described in Section 7.1, prescriptive DM/ML solutions are still quite application-bespoke and in need of generalization, e.g., as part of a tool, in order to be applicable to a broader set of different, but similar problems. As for prescriptive EXP workflows, domain-specific tools offering strong modelling and reasoning capabilities at a high level of abstraction might be a viable path forward.
@@
How to best offer tool support for emerging DPSA workflow patterns, be it in one unified tool, in several tools with good mutual integration, or otherwise, we leave as a possible research direction:
@@
RD4: Develop flexible and efficient tools for DPSA workflow patterns relying on prescriptive DM/ML or EXP.
10.5 Solutions for managing DPSA in production
In our selection criteria of this survey, we require under C5 that applications must be ”end-to-end” - i.e., that they must solve concrete problems with concrete methods and not just present hypotheticals (cf. Section 5.3). A stricter notion of ”end-to-end” concerns the entire life cycle of an application, including how the solution is deployed, monitored, and maintained in a production setting under changing requirements. One would expect that managing DPSA solutions in the wild is non-trivial. To mention a few example possibilities: They might need to interface with real-time control systems and/or people under strict time constraints; they might need safeguards and fallback mechanisms to prevent infeasible, dangerous, or untimely prescriptions; they might need specialized logging, instrumentation, and metrics to detect any performance deviations over time.
We are regrettably only able to give hypothetical examples, since our surveyed papers barely provided any information on how their solutions were, or might have been, deployed in a production setting, beyond simply stating that the solution was, will or could be ”deployed” (e.g., [85, 54, 59]). Other papers only consider individual issues that relate to the broader application life cycle - for example, the design of a dashboard to interface with their solution (e.g., [52, 126]), or the automation of their ML pipeline (e.g., [129]). Still, deployment details are surprisingly absent in our surveyed papers.
MLOps concerns how to properly deploy, monitor, and maintain ML solutions in production [175], which has been a major roadblock for the adoption of ML in industry [176]. Within the DPSA field, a narrow data scientist’s focus on how to combine various methods into novel workflows would leave it blind to important engineering issues related to the life cycle of its solutions, preventing them from broader adoption in industry. We therefore find that DPSA could benefit from exploring engineering-based approaches to ensure practical, maintainable, and robust deployments of its solutions - i.e., ”DecisionOps”, in a similar vein to MLOps. More concretely, we propose as a research direction to:
RD5: Develop novel architectures, methods, and tools to deploy, monitor, and maintain DPSA solutions in production.
@@ 11 Conclusion
In answering our three research questions in Sections 6-10, we have provided new conceptual models and temporal analyses of the field as a whole, both in terms of problem types (cf. Section 6), method types (cf. Sections 7), and workflow patterns (cf. Section 8). Based on temporal trends and pervasive Gordian Knots found within the set of surveyed papers, we have furthermore provided a foundation for future work by proposing a number of possible research directions (cf. RD1-5 from Section 10), both in terms of new methods, tools, and use cases. We can mnemonically summarize the key contributions in our conceptual framework as 10 applications domains, 5 method types, 2 generic workflow patterns, and 5 research directions. These contributions form a comprehensive model of the current state of the DPSA application research field and where it might be going in the future.